OpenStudy (anonymous):

for the givin proportion, find the unknown number n. 9/10 = n/6

OpenStudy (anonymous):

27/5

jhonyy9 (jhonyy9):

9/10 =n/6 54=10n n=54/10 =27/5 n=5,4

jhonyy9 (jhonyy9):

do you understand ?

jhonyy9 (jhonyy9):

it

OpenStudy (anonymous):

if n is 5 or 4 then how it's satisfying the equation

jhonyy9 (jhonyy9):

check the denominator common

jhonyy9 (jhonyy9):

no n not is 5 and 4 n=54/10

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

?

jhonyy9 (jhonyy9):

tankie put the good answer please if i have helped to you in this exercise

jhonyy9 (jhonyy9):

thank you

jhonyy9 (jhonyy9):

ta

jhonyy9 (jhonyy9):

tankie you have understood it all ?

jhonyy9 (jhonyy9):

ok what is your question ?

OpenStudy (anonymous):

see in unanswered list

jhonyy9 (jhonyy9):

ok i check it

OpenStudy (anonymous):

thank you

jhonyy9 (jhonyy9):

when you have wrote it ?

OpenStudy (anonymous):

two days have passed

jhonyy9 (jhonyy9):

till 14 hours ago i not have got it with your name nothing

OpenStudy (anonymous):

Can any body show me how?? [integrating from 0 to t]{f(t)*g(t-@)} d@=[integrating from 0 to infnity] {f(t)*g(t-@)*u(t-@) }d@ Here u(t) is the unit step function this was the question